# Autonomous Vehicles Using Multi-Agent Reinforcement Learning for Routing Decisions Can Harm Urban Traffic

### v1

@@ Abstract
Autonomous vehicles (AVs) using Multi-Agent Reinforcement Learning (MARL) for simultaneous route optimization may destabilize traffic environments, with human drivers possibly experiencing longer travel times. We study this interaction by simulating human drivers and AVs. Our experiments with standard MARL algorithms reveal that, even in trivial cases, policies often fail to converge to an optimal solution or require long training periods. The problem is amplified by the fact that we cannot rely entirely on simulated training, as there are no accurate models of human routing behavior. At the same time, real-world training in cities risks destabilizing urban traffic systems, increasing externalities, such as $CO_{2}$ emissions, and introducing non-stationarity as human drivers adapt unpredictably to AV behaviors. Centralization can improve convergence in some cases, however, it raises privacy concerns for the travelers’ destination data. In this position paper, we argue that future research must prioritize realistic benchmarks, cautious deployment strategies, and tools for monitoring and regulating AV routing behaviors to ensure sustainable and equitable urban mobility systems.
Multi-agent reinforcement learning, Route choice, Machine Learning, ICML
@@ 1 Introduction
Figure 1: Overview. We demonstrate how MARL fails to find optimal routing policies on a simple topology of two routes: shorter with no priority and longer with priority (a). We simulate 22 human drivers who reach two equilibria: (b) optimal, when all drivers use the shorter route 0, (c) and suboptimal, when all use the longer route 1. If we replace a single human driver with an AV, any standard RL algorithm will quickly find the optimal routing policy (d). However, when multiple AVs learn simultaneously, many MARL algorithms fail to converge to the optimal policy (e) or require hundreds of days (episodes) until they find it (f), even for such a trivial case.
@@
In urban traffic networks, humans (drivers) every day make routing decisions (Arnott et al., 1990) to arrive at their destinations as fast as possible (Bovy & Hoogendoorn-Lanser, 2005). With the advent of autonomous vehicles (AVs), these routing decisions may be delegated to algorithms, aiming to maximize the reward by selecting the optimal action in the current state of the network. Classically, this problem was formulated as a game-theoretical problem (Correa et al., 2004), where humans independently maximize their perceived payoffs.
Traditional methods that address the route choice problem often struggle to adapt to the complex and dynamic nature of mixed traffic scenarios involving AVs and human drivers (Bamdad Mehrabani et al., 2024). A plausible extension is to employ ML-equipped AVs and specifically use reinforcement learning, where each agent (vehicle) learns optimal policies to select the best route in the currently observed state of the urban road network. AVs can leverage approaches such as multi-agent reinforcement learning (MARL) to optimize their routing decisions for faster arrivals, benefiting fleet operators or individual clients. MARL has been used to solve this problem (Ramos et al., 2018; Thomasini et al., 2023; Tumer & Agogino, 2006), but from different perspectives, as discussed later.
@@
As AVs will start sharing the roads with humans in a mixed system, at least for a while until human driving ceases, they will be influencing the complex social dynamics of individual, rational yet non-deterministic driving behavior (Rahmati et al., 2019). However, as we argue machine intelligence will not have (as of today) a sufficiently detailed model of urban mobility capable of training routing algorithms. Furthermore, as we demonstrate, specifically during training, the joint actions of AVs may lead to suboptimal solutions, resulting in a cost to all users in the capacitated system with limited resources. Moreover, AVs are likely to take different actions than human drivers, which will likely trigger adaptations of humans. Humans then will change their routing policies in response to AVs’ actions. This impact is up to some level negligible and single or few AVs routing recklessly during training will not disequilibrate the system. Yet the critical mass of AVs can be quickly reached as the AVs become broadly available (as little as 15% of AVs in our experiment disequilibrate the system, Figure 5). There exist projections that by 2030, the number of commercial AVs for ridesharing will reach a few million (Sachs, 2024) and Statista (2024) projects that the number of AVs will rise significantly, from 17,000 in 2022 to 126,000 by 2030.
This position paper argues that even in trivial cases, multiple AVs simultaneously learning optimal MARL routing policies will either destabilize the road networks and fail to find optimal solutions or learn long enough to affect the system’s performance. We support this position with experiments on a toy network (Figure 1a) with a portfolio of MARL algorithms, discuss and show that:
  1. 1.
@@
    Human drivers, by maximizing their payoffs, stabilize the system into two equilibria (Figure 1b, 1c).
  2. 2.
@@
    RL efficiently finds the optimal route for a single AV (Figure 3). However, with multiple human drivers replaced by AVs, MARL either fails to converge or needs lengthy training to find the optimal solution (Figure 4).
  3. 3.
@@
    The simulators of urban mobility are not ready to serve as virtual environments to train MARL, and training in the real-world will be at the cost of the system’s performance (Section 2.4).
  4. 4.
@@
    Optimal equilibrium state can easily transition to suboptimal (with worse performance for each agent) when human adaptation is added as another source of non-stationarity (Figure 6 top).
  5. 5.
@@
    Centralization can, in some cases, accelerate the convergence to optimal policies, but at the cost of our privacy (Figure 6 bottom).
@@
This hinders the massive potential of AVs to contribute to sustainability (Taiebat et al., 2018), efficiency (Talebpour & Mahmassani, 2016), and optimality (Zhou et al., 2024a) of urban mobility. Connected and Autonomous Vehicles (CAVs) offer potential as they promise novel routing strategies that allow the reduction of total and individual costs (travel times) and system costs (like total delays) (Jamróz et al., 2024) and its externalities ($\text{CO}_{\text{2}}$, $\text{NO}_{\text{x}}$, safety, noise, etc.) (Kopelias et al., 2020).
@@
Based on findings from our illustrative experiments, to safely and reliably exploit the opportunities that AVs and MARL offer to the future urban traffic systems, we call for:
  1. 1.
    Broad experimental studies demonstrating the potential impact of MARL-enabled AVs on urban systems.
  2. 2.
    Benchmarks to test proposed MARL solutions in the standard set of networks, demand patterns, rich metrics, and efficient, rapidly converging MARL algorithms to simulate AVs’ decision-making.
  3. 3.
@@
    Data-driven development of urban traffic simulators, realistically reproducing human route choice (in the presence of real-time information) and its adaptation to dynamic changes in the environment.
  4. 4.
    Careful deployment of MARL routing systems in future AVs.
@@ 2 Background
2.1 Related work
MARL has been applied in many real-world applications such as finance, transportation, manufacturing (Zhou et al., 2024b) as well as chemistry and biology (Ning & Xie, 2024). For instance, Mi et al. (2024) introduced a simulation environment to model dynamic interactions among governments, households, firms, and financial intermediaries. Ma & Wu (2020) utilized a cooperative MARL approach to control the traffic signals in different road networks.
MARL has also been used to solve the route choice problems considering different aspects of the problem. Tumer & Agogino (2006) explored Q-learning for route selection, emphasizing reward-shaping techniques aiming to eliminate traffic congestion. Another study approached route choice as a congestion game aiming for a collective equilibrium in the traffic network (Zhou et al., 2020). Ramos et al. (2018) proposed a regret-minimization approach (Blum & Mansour, 2007) that relies on external traffic data, which may limit its applicability in scenarios where such information is incomplete or inaccurate. In a different vein, Thomasini et al. (2023) addressed route choice in a centralized multi-agent setting using a macroscopic traffic simulation. Lazar et al. (2021) address the route assignment problem to develop policies for CAVs that prevent unbounded delays in traffic networks. Akman et al. (2024) introduced AV-specific behavioral reward formulations in mixed-traffic environments.
@@ 2.2 Multi-agent reinforcement learning
The route assignment problem is a combinatorial optimization problem. A natural approach to this problem is to leverage an ML method capable of adapting to evolving and dynamic environments, such as urban networks. MARL is well-suited for representing AVs in traffic as it enables decentralized decision-making, adaptability to dynamic environments, and scalable multi-agent coordination.
@@
MARL involves multiple agents interacting within a shared environment, where each agent’s actions can influence the state of the environment. This makes the environment non-stationary from a single agent’s perspective. Employing a single-agent RL algorithm to learn value functions of joint actions would eliminate the non-stationarity (Claus & Boutilier, 1998; Hurtado et al., 2018). However, such a training setting does not scale well when the size of the action space grows exponentially with the number of agents (Lu et al., 2024) and requires a coordination mechanism.
@@
It is also possible to train a set of independent learners (IL) where each learner treats every other agent as a part of the environment. Our experiments start with Independent Q-Learning (IQL) as the initial baseline (Tan, 1997), followed by actor-critic methods such as the Independent SAC (ISAC), which is the multi-agent version of the Soft Actor-Critic (SAC) algorithm (Haarnoja et al., 2018a) and the Independent Proximal Policy Optimization (IPPO) algorithm, as it has demonstrated strong benchmark performance across a variety of problems (Yu et al., 2022; Papoudakis et al., 2021).
Existing literature widely utilizes the Centralized Training and Decentralized Execution (CTDE) structure, in which agents learn decentralized policies in a centralized manner, to be used independently during the execution phase (Lowe et al., 2020). This structure tackles partial observability, while maintaining feasible compute times.
@@
Following this structure, Sunehag et al. (2017) introduced Value Decomposition Networks (VDN), a value-based algorithm that decomposes complex learning problems into sub-problems. Rashid et al. (2018) proposed a similar algorithm named QMIX, which instead adopts a mixing network with a monotonicity constraint, enabling more complex coordination. Alternative to these value decomposition methods, Yu et al. (2021) introduced a multi-agent version of the Proximal Policy Optimization (PPO) algorithm (Schulman et al., 2017), Multi-Agent PPO (MAPPO). Similarly, Multi-Agent SAC (MASAC) represents the multi-agent version of the SAC algorithm (Haarnoja et al., 2018a). All these algorithms were employed to demonstrate our position in the experiments.
2.3 Partially Observable Markov Game (POMG)
We abstract the daily (repeated) route choices made by humans (and AVs in the future) in capacitated networks (limited resources) to the so-called routing game. This, in turn, can be formalized with a POMG as a MARL problem. We denote a finite-horizon multi-agent general-sum POMG as a tuple $(\mathcal{S},\{\mathcal{O}_{i}\}_{i\in\mathcal{N}},\{\mathcal{A}_{i}\}_{i\in\mathcal{N}},\mathcal{P},\{r_{i}\}_{i\in\mathcal{N}},\gamma)$, where $\mathcal{N}=\{0,\ldots,N-1\}$ represents the set of N agents. $\mathcal{S}$ represents a finite state space, partially observable by the agents. The global state $s\in\mathcal{S}$ includes the joint agent observation (i.e., $o_{1}\times o_{2}\times...\times o_{N}$) and additional environmental information. Each agent $i$ can apply an action $a_{i}^{t}\in\mathcal{A}_{i}$ at each timestep $t$ from its own action space $\mathcal{A}_{i}\subseteq\mathcal{A}$ to the environment and consequently receive a reward $r_{i}^{t}$ from the environment as a function of the state and the joint action. The state transition probability is defined by $\mathcal{P}(s^{\prime}\mid s,a):\mathcal{S}\times\mathcal{S}\times\mathcal{A}\to[0,1]$ and indicates the probability of transitioning to state $s^{\prime}\in\mathcal{S}$ after a joint action $a=(a_{1},a_{2},\dots,a_{N})\in\mathcal{A}$. A policy $\pi_{i}:\mathcal{O}_{i}\times\mathcal{A}_{i}\to[0,1]$ specifies the probability distribution over the actions of agent $i$. The goal of each agent $i$ is to learn a $\pi_{i}$ that maximizes its expected cumulative reward $G=\mathbb{E}_{\pi_{i}}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot r_{i}^{t}\right]$, where $\gamma\in[0,1)$ is the discount factor.
2.4 Are traffic models ready to train RL algorithms?
@@
    Traffic flow. We use Simulation of Urban MObility (SUMO) (Lopez et al., 2018), which applies Intelligent Driver Model (IDM) (Treiber et al., 2000) accurate, but not perfect microscopic model of traffic. Real traffic is less predictable and noisy and admits rare events like accidents (Chen et al., 2018).
  2. 2.
@@
    Demand patterns. In this paper, we assume a fixed commute pattern every day. Namely, each agent has a fixed origin from which it departs every day at the same time to a fixed destination. However, in real systems, humans change departure times, work irregularly, and occasionally stay at home or travel to different destinations (González et al., 2008; Horni et al., 2016; Bhat & Koppelman, 1999).
  3. 3.
@@
    Route choices. Humans are non-deterministic decision makers, and their choices are probabilistic with a significant impact of variation and heterogeneity on the choice probabilities (like in the Logit model (Ben-Akiva & Bierlaire, 1999)).
  4. 4.
@@
    Action space. In real networks, the number of feasible routes explodes quickly, reaching $10^{40}$ in many real-world examples (Frejinger et al., 2009).
@@
Unfortunately, all of the above are only roughly approximated with state-of-the-art transport system models. Our simple case involves 22 vehicles, two routes, and two equilibria that can be derived analytically. Although this setup is simple, it already gives rise to issues (like increased travel time for humans), and one can only expect them to worsen as the system scales to bigger networks with irregular demand patterns and non-deterministic traffic flows, which cannot be presented with state-of-the-art simulation models of urban mobility.
3 Methodology
We illustrate our position on a simple two-route network (Figure 1a), which seems abstract but can be identified in many complex urban topologies in the world (see Appendix A for examples). Each agent selects from two possible routes (0 or 1) to reach the destination. Then, the reward is collected from the environment to update choices for the next episode. Each episode is interpreted as a day on which each of the 22 drivers commutes through our network. For clarity, the setting is unrealistically static. Each human driver follows the same route every day, departing at the same time, and the travel times do not vary day-to-day (none of the above holds in real networks as discussed in Section 2.4). By design, the system has two Nash equilibria achievable by humans, one optimal (System Optimal - User Equilibrium, Figure 1b) and one suboptimal (Suboptimal - User equilibrium, Figure 1c). Humans mutate into AVs, which will use any MARL algorithms to find optimal policies. To simulate this scenario RouteRL (Psarou et al., 2024) environment is employed (Figure 2), which models mixed-traffic scenarios, where AVs are RL-enabled agents and human agents behave according to a given human-behavior model.
SUMO.
SUMO (Lopez et al., 2018) is a state-of-the-art, microscopic, agent-based traffic simulator, in which each vehicle navigates the road network according to the IDM.
Action.
The action space of the agents corresponds to the set of available routes connecting their given origin with their destination and is, thus, discrete with value two, on this network (See Figure 1a).
Reward.
The reward $r_{i}^{t}(a_{i}^{t}|s_{i}^{t})$ is the negative travel time of each agent $i$ to reach from its origin to its destination, as calculated by SUMO.
Observation.
We assume, plausibly for the future systems, that AVs observation $o_{i}^{t}$ is composed of their departure time and the number of agents that departed before them and chose routes 0 and 1, respectively.
Human agents.
We follow the classical representation of human route-choice behavior from transportation research. Human drivers are rational decision-makers aiming to maximize their perceived utility (Cascetta, 2009) by selecting actions that minimize expected travel times. Their expectations are updated based on experienced travel times (from SUMO). In scenarios with adaptation, humans shift to an alternative route with $10\%$ probability.

### v2

# POSITION: Collaboration Between the City and Machine Learning Community is Crucial to Efficient Autonomous Vehicles Routing

@@ Abstract
Autonomous vehicles (AVs), possibly using Multi-Agent Reinforcement Learning (MARL) for simultaneous route optimization, may destabilize traffic networks, with human drivers potentially experiencing longer travel times. We study this interaction by simulating human drivers and AVs. Our experiments with standard MARL algorithms reveal that, both in simplified and complex networks, policies often fail to converge to an optimal solution or require long training periods. This problem is amplified by the fact that we cannot rely entirely on simulated training, as there are no accurate models of human routing behavior. At the same time, real-world training in cities risks destabilizing urban traffic systems, increasing externalities, such as $CO_{2}$ emissions, and introducing non-stationarity as human drivers will adapt unpredictably to AV behaviors. In this position paper, we argue that city authorities must collaborate with the ML community to monitor and critically evaluate the routing algorithms proposed by car companies toward fair and system-efficient routing algorithms and regulatory standards.
@@ 1 Introduction
Figure 1: Overview. We demonstrate how MARL fails to find optimal routing policies on a simple topology of Two Routes: shorter with no priority (Yield sign) and longer with priority (TRY network). We simulate 22 human drivers who reach two equilibria: (a) optimal when all drivers use the shorter route 0, and (b) suboptimal when all use the longer route 1 (the first three humans always choose route 1 in this solution). If we replace a single human driver with an AV, any standard RL algorithm will quickly find the optimal routing policy (c). However, when multiple AVs (10) learn simultaneously, many MARL algorithms fail to converge to the optimal policy (d) or require hundreds of days (episodes) until they find it (e) in this trivial case as well as in real complex systems (Ingolstadt).
@@
In urban traffic networks, humans (drivers) every day make routing decisions arnott1990departure to arrive at their destinations as fast as possible bovy2005modelling . With the advent of autonomous vehicles (AVs), these routing decisions may be delegated to algorithms, aiming to maximize the reward by selecting the optimal action in the current state of the network. Classically, this problem was formulated as a game-theoretical problem correa2004selfish , where humans independently maximize their perceived payoffs.
@@
As AVs will start sharing the roads with humans in a mixed system, at least for a while until human driving ceases, they will be influencing the complex social dynamics of individual, rational yet non-deterministic human driving behavior influence_of_AVs_on_car_following_behavior . However, as we argue, current machine learning (ML) systems will not have (as of today) a sufficiently detailed model of urban mobility capable of training routing algorithms. Furthermore, as we demonstrate, specifically during training, the joint actions of AVs may lead to suboptimal solutions, resulting in a cost to all users in the capacitated system with limited resources. Moreover, AVs are likely to take different actions than human drivers, which will likely trigger adaptations of humans that will change their routing policies in response to AVs’ actions. This impact is negligible to some level, and single or few AVs routing recklessly during training will not disequilibrate the system. Yet the critical mass of AVs can be quickly reached as the AVs become broadly available (as little as 15% of AVs in our experiment disequilibrate the system). This aligns with projections indicating rapid growth in AV deployment, with the number of commercial AVs for ridesharing expected to reach a few million by 2030 goldman2024partiallyautonomous .
This position paper argues that city authorities must begin actively collaborating with the ML research community to monitor and evaluate AI-based routing algorithms for AVs deployed by car companies. In parallel, the ML community should focus on continuously improving existing algorithms to ensure they are robust, fair, and suitable for real-world deployment. Without regulation, autonomous collective routing could introduce chaos into urban traffic systems and exacerbate congestion. We support this position by demonstrating that, in simplified and real-world networks, when multiple AVs simultaneously learn routing policies using MARL, they will either destabilize the road networks and fail to find optimal solutions or learn long enough to affect the system’s performance. While MARL is not the only viable approach, it is the one we focus on in this study. Our argument is supported by experiments conducted on a toy network (Figure 1a, b) and complemented with experiments on the real-world Ingolstadt network (Figure 7), using a portfolio of MARL algorithms. We analyze and show that:
  • $\bullet$
@@
    Human drivers, by maximizing their payoffs, stabilize the system into two equilibria (Figure 1a, b).
  • $\bullet$
@@
    RL efficiently finds the optimal route for a single AV (Figure 3). However, when multiple human drivers are replaced by AVs, MARL either fails to converge or needs lengthy training to find the optimal solution (Figure 4(a, b)).
  • $\bullet$
@@
    The simulators of urban mobility are not ready to serve as virtual environments to train MARL, and training in the real world will be at the cost of the system’s performance (Section 2.3).
  • $\bullet$
@@
    Optimal equilibrium state can easily transition to suboptimal (with worse performance for each agent) when non-determinism (e.g., in the traffic simulation) and/or human adaptation is introduced as another source of non-stationarity (Figure 4c).
  • $\bullet$
@@
    Centralization can, in some cases, speed up convergence to optimal policies, but at the cost of our privacy (Figure 9 bottom).
  • $\bullet$
    In the real-world Ingolstadt network, multiple simultaneous routing AVs can destabilize the system (Figure 5).
@@
These phenomena can hinder the massive potential of AVs to contribute to sustainability taiebat2018review , efficiency talebpour2016influence , and optimality zhou2024modeling of urban mobility. While Connected and Autonomous Vehicles (CAVs) promise novel routing strategies that allow the reduction of total and individual costs (travel times) and system costs (like total delays) jamróz2024socialimpactcavs and its externalities ($\text{CO}_{\text{2}}$, $\text{NO}_{\text{x}}$, safety, noise, etc.) kopelias2020connected - these benefits can be limited by the challenges described above.
@@
Based on experiment findings, to safely and reliably exploit the opportunities that AVs and MARL offer to the future urban traffic systems, we call for:
  • $\bullet$
    The introduction of a regulatory framework, developed in collaboration with the ML community, that requires car companies to submit their routing algorithms for certification before deployment in public road networks.
  • $\bullet$
    The development and deployment of monitoring systems that track collective routing behaviors and can allow authorities to detect issues, like inequitable travel time distributions or prolonged system-wide congestion.
  • $\bullet$
@@
    The data-driven development of urban traffic simulators, led jointly by ML and transportation research, to realistically reproduce human route choice (in the presence of real-time information) and its adaptation to dynamic environmental changes.
  • $\bullet$
    The continuous development and improvement of ML algorithms that can robustly handle non-stationarity, scale with agent populations, and adapt to real-world traffic conditions.
  • $\bullet$
    Broad experimental studies to benchmark routing algorithms across diverse traffic scenarios and identify failure modes that inform regulatory policies and guide safe algorithm design.
@@ 2 Background
2.1 Partially Observable Markov Game (POMG)
We abstract the daily (repeated) route choices made by humans (and AVs in the future) in capacitated networks (limited resources) to the so-called routing game. This, in turn, can be formalized with a POMG as a MARL problem. We denote a finite-horizon multi-agent general-sum POMG as a tuple $(\mathcal{S},\{\mathcal{O}_{i}\}_{i\in\mathcal{N}},\{\mathcal{A}_{i}\}_{i\in\mathcal{N}},\mathcal{P},\{r_{i}\}_{i\in\mathcal{N}},\gamma)$, where $\mathcal{N}=\{0,\ldots,N-1\}$ represents the set of $N$ agents. $\mathcal{S}$ represents a finite state space, partially observable by the agents. The global state $s\in\mathcal{S}$ includes the joint agent observation (i.e., $o_{1}\times o_{2}\times...\times o_{N}$) and additional environmental information. Each agent $i$ applies an action $a_{i}^{t}\in\mathcal{A}_{i}$ at each timestep $t$ from its own action space $\mathcal{A}_{i}\subseteq\mathcal{A}$ to the environment and consequently receives a reward $r_{i}^{t}$ from the environment as a function of the state and the joint action. The state transition probability is defined by $\mathcal{P}(s^{\prime}\mid s,a):\mathcal{S}\times\mathcal{S}\times\mathcal{A}\to[0,1]$ and indicates the probability of transitioning to state $s^{\prime}\in\mathcal{S}$ after a joint action $a=(a_{1},a_{2},\dots,a_{N})\in\mathcal{A}$. A policy $\pi_{i}:\mathcal{O}_{i}\times\mathcal{A}_{i}\to[0,1]$ specifies the probability distribution over the actions of agent $i$. The goal of each agent $i$ is to learn a policy, $\pi_{i}$, that maximizes its expected cumulative reward $G=\mathbb{E}_{\pi_{i}}\left[\sum_{t=0}^{\infty}\gamma^{t}\cdot r_{i}^{t}\right]$, where $\gamma\in[0,1)$ is the discount factor.
@@ 2.2 Multi-agent reinforcement learning
The route assignment is a combinatorial optimization problem. A plausible extension is to employ ML-equipped AVs and specifically use multi-agent reinforcement learning, where each agent (vehicle) learns optimal policies to select the best route in the currently observed state of the road network. MARL involves multiple agents interacting within a shared environment, where each agent’s actions can influence the state of the environment. This makes the environment non-stationary from a single agent’s perspective. Using a single-agent RL algorithm to learn value functions of joint actions would eliminate the non-stationarity centralization ; ejal but would not scale well when the size of the action space grows exponentially with the number of agents lu2024centralized .
@@
One approach is to train a set of independent learners (IL) where each learner treats other agents as part of the environment. We use Independent Q-Learning (IQL) as the initial baseline iql , followed by two actor-critic methods, that is the Independent SAC (ISAC), the multi-agent version of the Soft Actor-Critic (SAC) algorithm sac , and the Independent Proximal Policy Optimization (IPPO) algorithm, as it has demonstrated benchmark performance across a variety of problems yu2022surprisingeffectivenessppocooperative ; papoudakis2021benchmarkingmultiagentdeepreinforcement .
@@
Existing literature widely utilizes the Centralized Training and Decentralized Execution (CTDE) structure, in which agents learn decentralized policies in a centralized manner, to be used independently during the execution phase lowe2020multiagentactorcriticmixedcooperativecompetitive . This structure tackles partial observability while maintaining feasible computation times. Following this structure, vdn introduced Value Decomposition Networks (VDN), a value-based algorithm that decomposes complex learning problems into sub-problems. qmix proposed a similar algorithm named QMIX, which adopts a mixing network with a monotonicity constraint, enabling more complex coordination. Alternative to value decomposition methods, mappo introduced a multi-agent version of the Proximal Policy Optimization (PPO) algorithm ppo , Multi-Agent PPO (MAPPO). Similarly, Multi-Agent SAC (MASAC) is the multi-agent version of the SAC algorithm. All these algorithms were used to demonstrate our position in the experiments.
2.3 Limitations: Are traffic models ready to train RL algorithms?
@@
    Traffic flow. We use Simulation of Urban MObility (SUMO) SUMO2018 , which applies Intelligent Driver Model (IDM) Treiber_2000 accurate, but not perfect microscopic model of traffic. Real traffic is less predictable and noisy and admits rare events like accidents chen2018exploring .
  • $\bullet$
@@
    Demand patterns. In this paper (except from Section 5.6), we assume a fixed commute pattern every day. Namely, each agent has a fixed origin from which it departs every day at the same time to a fixed destination. However, in real systems, humans change departure times, work irregularly, and occasionally stay at home or travel to different destinations Gonz_lez_2008 ; horni2016matsim ; Bhat1999 .
  • $\bullet$
@@
    Route choices. Humans are non-deterministic decision makers, and their choices are probabilistic with a significant impact of variation and heterogeneity on the choice probabilities (like in the Logit model ben1999discrete ).
  • $\bullet$
@@
    Action space. In real networks, the number of feasible routes explodes quickly, reaching $10^{40}$ in many real-world examples frejinger2009sampling .
@@
Unfortunately, all of the above are only roughly approximated with state-of-the-art transport system models. Although the setup discussed in Figure 1 is simple, it already gives rise to issues (like increased travel time for humans), and these worsen as the system scales to bigger networks (replicated using the Ingolstadt network) with irregular demand patterns and non-deterministic traffic flows (see Figure 4c). In these cases, we get similarly bad results: MARL fails, supporting our position.
3 Related work
3.1 Analytical solutions
While the nature of the problem and its formalization as a POMG renders (MA)RL likely to be the tool of choice to solve fleet route choice problems we do not argue that it is the best and only solution. The classic traffic assignment solvers identify user equilibrium and system optimum using Operations Research (OR) methods, like Frank-Wolfe (fukushima1984modified, ). However, these methods operate on macroscopic flows and convex, continuous travel-time functions kucharski2017estimating , while we study an agent-based setting with a realistic, microscopic traffic flow model (SUMO). Alternatively, game-theoretical equilibria solvers, like Gambit turocy2001gambit , can be used. However, Gambit requires a precomputed payoff matrix, that includes $2^{10}$ episode runs to account for all the possible joint action combinations and the complexity of the game grows exponentially. Hence, such solvers often struggle to adapt to the complex and dynamic nature of mixed traffic scenarios involving AVs and human drivers BAMDADMEHRABANI2024 and scale poorly in bigger networks, highlighting the need for alternatives, with MARL emerging as the most promising.
3.2 Multi-agent reinforcement learning
RL has been applied to the route choice problem in a macroscopic setting, where vehicle groups are modeled collectively as flows instead of discrete agents. For example, Thomasini+2023 studied route choice in a centralized multi-agent setting using a macroscopic traffic simulator. Additionally, ZHOU2020124895 formulated route choice as a congestion game transforming it into the Traffic Assignment Problem (TAP). They proposed an RL-based solution that converges almost surely to the optimal solution aiming to minimize the total travel time in traffic networks. While macroscopic approaches explore overall traffic dynamics, they abstract away the heterogeneity and strategic behavior of individual agents.
In contrast, microscopic approaches focus on individual vehicle decisions, making them more suitable for modeling AV–human interactions. For instance, AgentRewardShaping adopted a multi-agent approach where drivers use Q-learning for route selection using reward-shaping techniques to reduce traffic congestion. Additionally, regret_route_choice proposed a regret-minimization approach that relies on external traffic data.
Some studies model route choice as a sequence of decisions made at each network node indi_vs_diff_rewards_route_choice ; improved_learning_automata_route_choice , allowing for dynamic adaptation to changing traffic conditions. In a different vein, lazar2021learningdynamicallyrouteautonomous consider a setting where human drivers act selfishly and AVs centrally controlled using deep RL decrease congestion by indirectly influencing human’s routing decisions. Lastly, akman2024impact introduced AV-specific behavioral reward formulations in mixed-traffic environments which are later included in the RouteRL framework RouteRL2024 for simulating the collective route choices of both human drivers and AVs.
4 Problem statement